OpenStudy (anonymous):

Evaluate 15 - 4 + (7 - 5)2

OpenStudy (anonymous):

15

OpenStudy (anonymous):

15 15-4+(7-5)x2 7-5=2 15-4+2X2 2x2=4 15-4+4 15-4=11 11+4= 15 15

OpenStudy (anonymous):

oops, it actually looks like this http://prntscr.com/38qn5c

OpenStudy (anonymous):

use wolframalpha.com/ to compute your things it always works for me.

OpenStudy (anonymous):

:O thanks @grabill

OpenStudy (anonymous):

your welcome :)